Suppose that in January 2006 Kenneth Cole Productions had EPS of $ 1.76 and a book value of equity of $ 14.32 per share. Use the multiples approach to estimate? KCP's value based on the data from comparable firms given in the following? table: LOADING....

a. Using the average? P/E multiple from the table? above, estimate? KCP's share price. (Round to the nearest cent)

b. What range of share prices do you estimate based on the highest and lowest? P/E multiples in the table? above? (Round to the nearest cent)

c. Using the average price to book value multiple in the table? above, estimate? KCP's share price. (Round to the nearest cent)

d. What range of share prices do you estimate based on the highest and lowest price to book value multiples in the table? below? Round to the nearest cent?

P/E Price/Book Enterprise Value/Sales enterprise Value/EBITDA
Average 15.01 2.84 1.06 8.49
Maximum +51% +186% +106% +27%
Minimum -42% -61% -56% -22%
Answer #1

As per rules I will answer the first 4 sub parts of this question

EPS=$1.76

Book value of equity= $14.32

1)Average P/E Ratio = 15.01 = Price/ EPS

Hence price= 15.01*1.76 = $26.42

2) Highest P/E = 15.01*151% =22.6651

Highest price= 22.6651*1.76 = $39.89

3) Lowest P/E = 15.01*(1-0.42)= 8.7058

Lowest Price = 8.7057*1.76 = $15.31

4)Average Price to book value = 2.84

Price of the share = Price/Book Ratio*Book value of equity

=2.84*14.32 = $40.67